Hi...my name is kendra and im reallly trying to pull up my grade in math...I dont really know when a problem is undefined and in seeming to always get problem like that wrong..can you please please help me
Answer by stanbon(75887)

Hi...my name is kendra and im reallly trying to pull up my grade in math...I dont really know when a problem is undefined and in seeming to always get problem like that wrong..can you please please help me
-----------------
Problems are frequently undefined because:
A variable is in a denominator and the problem
is undefined when that denominator is zero.
Example f(x) = 2/(x-3) is undefined for x=3.
-----------------
Ex: g(x) = 3/[x(x-2)] is undefined for x = 0
and for x = 2.
-----------------
Cheers,
Stan H.